# Complete $k$-partite entanglement measure

### v2

@@ IV.2 $k$-PEMo from unified MEM
In general, for any given pure state $|\psi\rangle=|\psi\rangle^{A_{1}A_{2}\cdots A_{n}}$ in $\mathcal{H}^{A_{1}A_{2}\cdots A_{n}}$, we assume it is not $(k-1)$-producible. Then there exists a $l$-fitness partition $X_{1}|X_{2}|\cdots|X_{m}$, $l\geqslant k$, such that
@@
for some subsystem $X_{t}$ in the partition $X_{1}|X_{2}|\cdots|X_{m}$. Let $t_{1}$, $t_{2}$, $\dots$, $t_{l}$ be all of the subscripts such that $X_{t_{i}}$ satisfies the condition (19) corresponding to all possible $l$-fitness partitions with $l\geqslant k$. It turns out that
$|\psi\rangle=|\psi\rangle^{X_{t_{1}}}|\psi\rangle^{X_{t_{2}}}\cdots|\psi\rangle^{X_{t_{l}}}|\phi\rangle^{X_{*}}$ (20)
under some permutation of the subsystems, where $X_{*}$ denotes the subsystem complementary to $X_{t_{1}}X_{t_{2}}\cdots X_{t_{l}}$. In such a sense, we can quantify the $k$-partite entanglement of $|\psi\rangle$ by
$\check{E}_{(k)}(|\psi\rangle)=\sum\limits_{j=1}^{l}E^{(s(t_{j}))}(|\psi\rangle^{X_{t_{j}}})$ (21)
@@
$\check{E}_{(k)}(|\psi\rangle)=\begin{cases}\sum\limits_{j=1}^{l}E^{(s(t_{j}))}(|\psi\rangle^{X_{t_{j}}}),&E_{(k)}(|\psi\rangle)>0,\\ 0,&E_{(k)}(|\psi\rangle)=0.\end{cases}\quad~{}$ (22)
@@
Let $h$ be a non-negative concave function. If $E^{(n)}(|\psi\rangle)=\frac{1}{2}\sum_{i}h(\rho^{A_{i}})$, then $\check{E}_{(k)}$ is a unified $k$-PEMo, and $\check{E}_{(2)}$ is a complete $2$-partite EMo whenever the reduced function is subadditve.
@@
Items (i) and (ii) are straightforward. For any $|\psi\rangle\in\mathcal{S}^{A_{1}A_{2}\cdots A_{n}}$, we suppose that $\check{E}_{(k)}(|\psi\rangle)=\sum\limits_{j=1}^{l}E^{(s(t_{j}))}(|\psi\rangle^{X_{t_{j}}})>0$ for some partition $X_{t_{1}}|X_{t_{2}}|\cdots|X_{t_{l}}|X_{*}$. It turns out that $\check{E}_{(k)}(|\psi\rangle^{X_{t_{j}}})=\check{E}_{(k-1)}(|\psi\rangle^{X_{t_{j}}})$ for any $j$, $\check{E}_{(k)}(|\phi\rangle^{X_{*}})=0$, but it is possible that $\check{E}_{(k-1)}(|\phi\rangle^{X_{*}})>0$, which implies (iii) is true. For any partition $X_{1}|X_{2}|\cdots|X_{p}$ of $A_{1}A_{2}\cdots A_{n}$, $p\geqslant k$, we consider $X_{2}|\cdots|X_{p}$ w.n.l.g., namely the partition that by discarding $X_{1}$ from $X_{1}|X_{2}|\cdots|X_{p}$. If $X_{1}$ is $X_{t_{j}}$ or some subsytem(s) of $X_{t_{j}}$, (iv) is clear since $E^{(s)}$ is unified [26] [a unified MEM is decreasing under the coarsening relation of type (a)]. If $X_{1}$ is $X_{*}$ or some subsytem(s) of $X_{*}$, (iv) is clear since $\check{E}_{(k)}(|\psi\rangle)=\check{E}_{(k)}(\rho^{X_{2}|\cdots|X_{p}})$. If $X_{1}$ cotains some subsytem(s) of $X_{t_{j}}$, (iv) is also ture since $E^{(s)}$ is unified. The other cases are obvious.
For any partition $X_{1}|X_{2}|\cdots|X_{p}$ of $A_{1}A_{2}\cdots A_{n}$, $p\geqslant k$, we consider $X_{1}X_{2}X_{3}|X_{4}|\cdots|X_{p}$ w.n.l.g., namely the partition that by combining $X_{1}$, $X_{2}$ and $X_{3}$ from $X_{1}|X_{2}|\cdots|X_{p}$. There are two different cases: (a) For any $X_{t_{j}}$, either $X_{t_{j}}$ is some subysystem of $X_{1}X_{2}X_{3}$, or $X_{t_{j}}$ is some subysystem of $X_{4}\cdots X_{p}$, (b) There exist some $X_{t_{j}}=A_{t_{j},1}A_{t_{j},2}\cdots A_{t_{j},s}$ such that some $A_{t_{j},i}$ is/are subsystem(s) of $X_{1}X_{2}X_{3}$ while $\overline{A_{t_{j},i}}$ is/are subsystem(s) of $X_{4}\cdots X_{p}$. The case of (a) is clear, and the case of (b) is also true since the reduced function is subadditive. ∎
However $\check{E}_{(k)}$ are not complete $k$-PEMos for any $k\geqslant 3$. For example, we take
then
$\check{E}_{(3)}(|\psi\rangle)=\check{E}_{(3)}(|\psi\rangle^{ABC}).$
@@
$\check{E}_{(3)}(A|B|C|D|EFGI|H|J)$
$=$ $\check{E}_{(3)}(ABC)+\check{E}_{(3)}(D|EFGI|H|J)$
which is larger than $\check{E}_{(3)}(|\psi\rangle)$ whenever $|\psi\rangle^{DE}$, $|\psi\rangle^{GH}$ and $|\psi\rangle^{IJ}$ are entangled states.
Another candidate for the unified global multipartite entanglement measure is the one defined by the sum of all bipartite entanglement [24], i.e.,
$\mathcal{E}^{(n)}(|\psi\rangle^{A_{1}A_{2}\cdots A_{n}})$ (25)
$=$ $\left\{\begin{array}[]{ll}\frac{1}{2}\sum\limits_{i_{1}\leq\cdots\leq i_{s},s<n/2}h(\rho^{A_{i_{1}}A_{i_{2}}\cdots A_{i_{s}}}),&\mbox{if}~{}n~{}\mbox{is odd},\\ \frac{1}{2}\sum\limits_{i_{1}\leq\cdots\leq i_{s}<n,s\leq n/2}h(\rho^{A_{i_{1}}A_{i_{2}}\cdots A_{i_{s}}}),&\mbox{if}~{}n~{}\mbox{is even},\end{array}\right.~{}~{}~{}~{}~{}$
where $h$ is a non-negative concave function. We denote by ˇ$\check{\mathcal{E}}_{(k)}$ the quantity that is defined as in Eq. (22) just with $\mathcal{E}^{(s(t_{j}))}$ replacing $E^{(s(t_{j}))}$. By definition $\check{E}_{g(2)}=E_{g(2)}$, and if the reduced function is subadditive, it can be easily checked that
$E_{(k)}(\rho)\leqslant\check{E}_{(k)}(\rho)\leqslant\check{\mathcal{E}}_{(k)}(\rho).$ (26)
Using similar arguments as in the proof of Theorem 2, we can conclude the following theorem.
Theorem 3.
ˇ$\check{\mathcal{E}}_{(k)}$ is a unified $k$-PEMo. ˇ$\check{\mathcal{E}}_{(2)}$ is a complete $k$-PEMo if the reduced function is subadditive while any $k$-PEM is not complete for $k\geq 3$.
In Eq. (22), $|\psi\rangle^{X_{t_{j}}}$ is genuinely entangled, so $E^{(s(t_{j}))}$ can choose any genuine entanglement measure instead. For example, the GMEM from the minimal reduced function $E_{g^{\prime\prime}}^{(n)}$ [26], which is defined by
$E_{g^{\prime\prime}}^{(n)}(|\psi\rangle^{A_{1}A_{2}\cdots A_{n}})=\min\limits_{i}h(\rho^{A_{i}}),$
where $h$ is a non-negative concave function. Then the corresponding $k$-PEMo is not unified in general since $E_{g^{\prime\prime}}^{(n)}$ may decrease under the coarsening relation of type (a) [26].
By definitions, both $E_{(k)}$ and $\check{E}_{(k)}$ are not genuine $k$-partite entanglement measures. If $E_{(k)}(|\psi\rangle)>0$ (resp. $\check{E}_{(k)}(|\psi\rangle)>0$) but $E_{(k+1)}(|\psi\rangle)=0$ (resp. $\check{E}_{(k+1)}(|\psi\rangle)=0$), then $|\psi\rangle$ is genuine $k$-partite entangled.
V Examples
We illustrate $E_{(k)}$, $\check{E}_{(k)}$, and ˇ$\check{\mathcal{E}}_{(k)}$ with the reduced functions $h_{C}(\rho)=\sqrt{2(1-{\rm Tr}\rho^{2})}$ and $h(\rho)=S(\rho)$ (i.e., $h_{C}$ is the reduced function of concurrence [45], $S$ is the reduced function of the entanglement of formation [46]), respectively. We denote them by $C_{(k)}$, $\check{C}_{(k)}$, and $\check{\mathcal{C}}_{(k)}$ if the reduced function is $h_{C}$, and by $E_{(k)}$, $\check{E}_{(k)}$, and ˇ$\check{\mathcal{E}}_{(k)}$ whenever the reduced function is $S$. Since $S$ and $h_{C}$ are subadditive [47, 26], so $E_{(2)}$, $\check{E}_{(2)}$, ˇ$\check{\mathcal{E}}_{(2)}$, $C_{(2)}$, $\check{C}_{(2)}$, and $\check{\mathcal{C}}_{(2)}$ are complete.

### v3

@@ 4.1 $k$-PEMo from unified multipartite entanglement measure
In general, for any given pure state $|\psi\rangle=|\psi\rangle^{A_{1}A_{2}\cdots A_{n}}$ in $\mathcal{H}^{A_{1}A_{2}\cdots A_{n}}$, we assume it is not (k1)(k-1)-producible. Then there exists a $l$-fitness partition $X_{1}|X_{2}|\cdots|X_{m}$, $l\geqslant k$, such that
@@
for some subsystem $X_{t}$ in the partition $X_{1}|X_{2}|\cdots|X_{m}$. Let $t_{1}$, $t_{2}$, $\dots$, $t_{r}$ be all of the subscripts such that $X_{t_{i}}$ satisfies the condition (17) corresponding to all possible $l$-fitness partitions with $l\geqslant k$. It turns out that
$|\psi\rangle=|\psi\rangle^{X_{t_{1}}}|\psi\rangle^{X_{t_{2}}}\cdots|\psi\rangle^{X_{t_{r}}}|\phi\rangle^{X_{*}}$ (18)
under some permutation of the subsystems, where $X_{*}$ denotes the subsystem complementary to $X_{t_{1}}X_{t_{2}}\cdots X_{t_{r}}$. In such a sense, we can quantify the $k$-partite entanglement of $|\psi\rangle$ by
$E_{(k)}(|\psi\rangle)=\sum\limits_{j=1}^{r}E^{(s(t_{j}))}(|\psi\rangle^{X_{t_{j}}})$ (19)
@@
$E_{(k)}(|\psi\rangle)=\begin{cases}\sum\limits_{j=1}^{r}E^{(s(t_{j}))}(|\psi\rangle^{X_{t_{j}}}),&s(t_{j})\geq k~\text{for some}~j,\\ 0,&\text{otherwise}.\end{cases}$ (20)
@@
Let $h$ be a reduced function. If $E^{(n)}(|\psi\rangle)=\frac{1}{2}\sum_{i}h(\rho^{A_{i}})$, then $E_{(k)}$ is a unified $k$-PEMo, and moreover, $E_{(k)}$ is a complete $k$-PEMo whenever $h$ is subadditve.
@@
Items (i) and (ii) are straightforward. For any $|\psi\rangle\in\mathcal{S}^{A_{1}A_{2}\cdots A_{n}}$, we suppose that $E_{(k)}(|\psi\rangle)=\sum\limits_{j=1}^{r}E^{(s(t_{j}))}(|\psi\rangle^{X_{t_{j}}})>0$ for some partition $X_{t_{1}}|X_{t_{2}}|\cdots|X_{t_{r}}|X_{*}$. It turns out that $E_{(k)}(|\psi\rangle^{X_{t_{j}}})=E_{(k-1)}(|\psi\rangle^{X_{t_{j}}})$ for any $j$, $E_{(k)}(|\phi\rangle^{X_{*}})=0$, but it is possible that $E_{(k-1)}(|\phi\rangle^{X_{*}})>0$, which implies (iii) is true. For any partition $X_{1}|X_{2}|\cdots|X_{p}$ of $A_{1}A_{2}\cdots A_{n}$, $p\geqslant k$, we consider $X_{2}|\cdots|X_{p}$ w.n.l.g., namely the partition that by discarding $X_{1}$ from $X_{1}|X_{2}|\cdots|X_{p}$. If $X_{1}$ is $X_{t_{j}}$ or some subsytem(ssome) of $X_{t_{j}}$, (iv) is clear since $E^{(s)}$ is unified [24] [a unified MEM is decreasing under the coarsening relation of type (a)]. If $X_{1}$ is $X_{*}$ or some subsytem(s) of $X_{*}$, (iv) is clear since $E_{(k)}(|\psi\rangle)=E_{(k)}(\rho^{X_{2}|\cdots|X_{p}})$. If $X_{1}$ cotains some subsytem(s) of some $X_{t_{j}}$, (iv) is also ture since $E^{(s)}$ is unified. The other cases can be argued similarly.
If the reduced function $h$ is subadditive, then $E^{(n)}$ is complete [21, 24], which means that $E^{(n)}$ is nonincreasing under the coarsening operation of type (b). This completes the proof. ∎
Another candidate for the unified global MEM is the one defined by the sum of all bipartite entanglement [22], i.e.,
$\qquad\mathcal{E}^{(n)}(|\psi\rangle^{A_{1}A_{2}\cdots A_{n}})=\left\{\begin{array}[]{ll}\frac{1}{2}\sum\limits_{i_{1}\leq\cdots\leq i_{s},s<n/2}h(\rho^{A_{i_{1}}A_{i_{2}}\cdots A_{i_{s}}}),&\mbox{if}~n~\mbox{is odd},\\ \frac{1}{2}\sum\limits_{i_{1}\leq\cdots\leq i_{s}<n,s\leq n/2}h(\rho^{A_{i_{1}}A_{i_{2}}\cdots A_{i_{s}}}),&\mbox{if}~n~\mbox{is even},\end{array}\right$.~~~~~ (23)
where $h$ is some given reduced function. We denote by $\mathcal{E}_{(k)}$ the quantity that is defined as in Eq. (20) just with $\mathcal{E}^{(s(t_{j}))}$ replacing $E^{(s(t_{j}))}$. Using similar arguments as in the proof of Theorem 1, we can conclude the following theorem.
Theorem 2.
$\mathcal{E}_{(k)}$ is a unified $k$-PEMo, and $\mathcal{E}_{(k)}$ is complete if $h$ is subadditive.
In Eq. (20), $|\psi\rangle^{X_{t_{j}}}$ is genuinely entangled, so $E^{(s(t_{j}))}$ can be any genuine entanglement measure instead. For example, the genuine entanglement measure from the minimal reduced function, which is defined by [24]
$\varepsilon_{g^{\prime\prime}}^{(n)}(|\psi\rangle^{A_{1}A_{2}\cdots A_{n}})=h(\rho^{X})$
where $h$ is some given reduced function, $X\subsetneq\{A_{1},A_{2},\dots,A_{n}\}$, namely the minimum runs over all possible reduced states. Then the corresponding $k$-PEMo is not unified in general since $\varepsilon_{g^{\prime\prime}}^{(n)}$ may increase under the coarsening relation of type (a) [24].
4.2 $k$-PEMo from the minimal sum
Let $|\psi\rangle=|\psi\rangle^{A_{1}A_{2}\cdots A_{n}}$ be a pure state in $\mathcal{H}^{A_{1}A_{2}\cdots A_{n}}$ and $h$ be a reduced function. For any $\gamma_{i}^{f}\in\Gamma_{k}^{f}$, we write
[…]
$E^{\prime}_{(k)}(|\psi\rangle)=\min\limits_{\Gamma_{\!k-\!1}^{f}}\mathcal{P}_{\!k-\!1}^{\gamma_{i}^{f}}(|\psi\rangle),$ (25)
where the minimum is taken over all feasible (k1)(k\!\!-\!\!1)-fineness partitions in $\Gamma_{\!k-\!1}^{f}$. By definition, for any $\rho\in\mathcal{S}^{A_{1}A_{2}\cdots A_{n}}$, $E^{\prime}_{(k)}(\rho)>0$ if and only if $\rho$ is $k$-partite entangled.
Theorem 3.
$E^{\prime}_{(k)}$ is a unified $k$-PEMo and $E^{\prime}_{(2)}$ is complete if the reduced function is subadditive.
Proof.
By definition, it is straightforward that $E^{\prime}_{(k)}$ is a $k$-PEMo. We show below it satisfies the unification conditions (i)-(iv). We only need to check items (iii) and (iv) since (i) and (ii) are clear. Since $\Gamma^{f}_{k\!-\!2}\subseteq\Gamma^{f}_{k\!-\!1}$, this implies (iii) is true. For any given $|\psi\rangle\in\mathcal{H}^{A_{1}A_{2}\cdots A_{n}}$, we assume w.n.l.g. that
$E^{\prime}_{(k)}(|\psi\rangle)=\frac{1}{2}\left[h(\rho^{X_{1}})+h(\rho^{X_{2}})+h(\rho^{X_{3}})\right]$
for some $X_{1}|X_{2}|\cdots|X_{m}$. If
$Y_{1}|Y_{2}|\cdots|Y_{p}\succ^{a}Y^{\prime}_{1}|Y^{\prime}_{2}|\cdots|Y^{\prime}_{q}$
with $Y^{\prime}_{1}|Y^{\prime}_{2}|\cdots|Y^{\prime}_{q}$ is obtained by discarding subsystem(s) $Y_{i}$s such that $Y_{i}$s are contained in $X_{4}X_{5}\cdots X_{m}$, Eq. (12) is clear. If $Y^{\prime}_{1}|Y^{\prime}_{2}|\cdots|Y^{\prime}_{q}$ is obtained by discarding subsystem(s) $Y_{i}$s such that $Y_{i}$s are contained in $X_{1}X_{2}X_{3}$, there are three subcases: (a) $Y_{i}=X_{1}$ (w.n.l.g.), (b) $Y_{i}=X_{21}$, $X_{2}=X_{21}X_{22}$ (w.n.l.g.), and (c) $Y_{i}=X_{1}X_{21}$ (w.n.l.g.). For the subcase of (a), it turns out that
$E^{\prime}_{(k)}(|\psi\rangle^{X_{1}X_{2}X_{3}})$
$=$ $\frac{1}{2}\left[h(\rho^{X_{1}})+h(\rho^{X_{2}})+h(\rho^{X_{3}})\right]$
$>$ $\frac{1}{2}\sum_{i}p_{i}\left[h(\rho^{X_{2(i)}})+h(\rho^{X_{3(i)}})\right]$
$\geqslant$ $E^{\prime}_{(k)}(\rho^{X_{2}X_{3}})$
for any pure state ensemble $\{p_{i},|\psi\rangle^{X_{2(i)}X_{3(i)}}\}$ of $\rho^{X_{2}X_{3}}$ since $h$ is concave. That is, Eq. (12) holds true still. For the subcase of (b), it follows that
$\quad E^{\prime}_{(k)}(|\psi\rangle^{X_{1}X_{2}X_{3}})\geqslant\frac{1}{2}\sum_{i}p_{i}\left[h(\rho^{X_{1(i)}})+h(\rho^{X_{22(i)}})+h(\rho^{X_{3(i)}})\right]\geqslant E^{\prime}_{(k)}(\rho^{X_{12}X_{2}X_{3}})$
for any pure state ensemble $\{p_{i},|\psi\rangle^{X_{1(i)}X_{22(i)}X_{3(i)}}\}$ of $\rho^{X_{1}X_{22}X_{3}}$, where the first inequality holds since discarding $X_{11}$ is a partial trace which is a special LOCC on $X_{1}|X_{2}|X_{3}$, and $E^{\prime}_{(k)}$ in such a case is $E^{(3)}$ acting on $\mathcal{S}^{X_{1}X_{2}X_{3}}$ which leads to decreasing under LOCC. The subcase of (c) is clear from the arguments for (a) together with (b). That is, Eq. (12) is valid for both subcases.
It is clear that the completeness of $E^{\prime}_{(2)}$ is reduced to the subadditivity of the reduced function. This completes the proof. ∎
By definition, if the reduced function is subadditive, it can be easily checked that
$E^{\prime}_{(k)}(\rho)\leqslant E_{(k)}(\rho)\leqslant\mathcal{E}_{(k)}(\rho).$ (26)
If $h$ is subadditive, then the minimal partition is the ones that contained in $\Gamma_{\!k-\!1}^{f}\!\!\setminus\!\Gamma_{\!k-\!2}^{f}$. $E^{\prime}_{(k)}$ is not complete in general if $k\geqslant 3$. For example, if
$\qquad\quad E^{\prime}_{(3)}(|\psi\rangle^{ABCD}|\psi\rangle^{EF}|\psi\rangle^{GHI})=\frac{1}{2}\left[h(\rho^{AB})+h(\rho^{CD})+h(\rho^{GH})+h(\rho^{I})\right],$
it follows that
$E^{\prime}_{(3)}(A|BC|D|E|F|G|H|I)$
$=$ $\frac{1}{2}\left[h(\rho^{A})+h(\rho^{BCD})+h(\rho^{GH})+h(\rho^{I})\right]$
$\geqslant$ $E^{\prime}_{(3)}(ABCDEFGH)$
whenever $h(\rho^{D})\geq h(\rho^{A})>h(\rho^{AB})$.
$\qquad\quad C_{(3)}(|\psi\rangle^{ABC}|\psi\rangle^{DE}|\psi\rangle^{FGH})=\frac{1}{5}\left[h(\rho^{AB})+h(\rho^{C})+h(\rho^{FG})+h(\rho^{H})\right],$
then $C_{(2)}(|\psi\rangle^{ABC}|\psi\rangle^{DE}|\psi\rangle^{FGH})=\frac{1}{8}(h_{A}+h_{B}+h_{C}+h_{D}+h_{E}+h_{F}+h_{G}+h_{H})$. Hereafter, we denote $h(\rho^{X})$ by $h_{X}$ for simplicity. Clearly, it is not necessary that $C_{(3)}(|\psi\rangle^{ABC}|\psi\rangle^{DE}|\psi\rangle^{FGH})\leq C_{(2)}(|\psi\rangle^{ABC}|\psi\rangle^{DE}|\psi\rangle^{FGH})$. So it is not $k$-monotonic. Let
$\qquad\quad C_{(3)}(|\psi\rangle^{AB}|\psi\rangle^{C}|\psi\rangle^{DEF})=\frac{1}{4}(h_{AB}+h_{C}+h_{DE}+h_{F})=\frac{1}{4}(h_{DE}+h_{F}).$
Then $C_{(3)}(|\psi\rangle^{AB}|\psi\rangle^{DEF})=\frac{1}{3}(h_{AB}+h_{DE}+h_{F})=\frac{1}{3}(h_{DE}+h_{F})$, i.e., $C_{(3)}$ is not coarsening monotonic. In addition $C_{(3)}(|\psi\rangle^{AB}|\psi\rangle^{C}+C_{(3)}(|\psi\rangle^{DEF})=C_{(3)}(|\psi\rangle^{DEF})=\frac{1}{2}(h_{DE}+h_{F})$, so $C_{(3)}$ is not additive. For $|\psi\rangle^{AB}|\psi\rangle^{CD}$, $C_{(2)}(|\psi\rangle^{AB}|\psi\rangle^{CD})=\frac{1}{4}(h_{A}+h_{B}+h_{C}+h_{D})=\frac{1}{2}(h_{A}+h_{C})$ may be not larger than $C_{(2)}(AB|C|D)=2h_{C}/3$.
Take $|\psi\rangle^{AB}|\psi\rangle^{C}|\psi\rangle^{DE}$, then
$C_{q(2)}(|\psi\rangle^{AB}|\psi\rangle^{C}|\psi\rangle^{DE})=\sqrt{\frac{2}{5}}\sqrt{h_{A}+h_{D}}$
$\neq$ $C_{q(2)}(|\psi\rangle^{AB})+C_{q(2)}(|\psi\rangle^{C}|\psi\rangle^{DE})=\sqrt{h_{A}}+\sqrt{2h_{D}/3},$
in general. So it is not additive. For $|\psi\rangle^{AB}|\psi\rangle^{C}|\psi\rangle^{DEF}$, we assume that
$C_{q(3)}(|\psi\rangle^{AB}|\psi\rangle^{C}|\psi\rangle^{DEF})=\sqrt{h_{DE}+h_{F}}/2.$
@@ 4.2 $k$-PEMo from the minimal sum
$C_{q(2)}(|\psi\rangle^{AB}|\psi\rangle^{C}|\psi\rangle^{DEF})=\sqrt{(h_{A}+h_{B}+h_{D}+h_{E}+h_{F})/6},$
which can not guarantee $C_{q(3)}\leq C_{q(2)}$. In addition,
$\qquad C_{q(3)}(|\psi\rangle^{AB}|\psi\rangle^{C}|\psi\rangle^{DEF})=\sqrt{h_{F}}/2<C_{q(3)}(|\psi\rangle^{AB}|\psi\rangle^{DEF})=\sqrt{2h_{F}/3}$
implies that $C_{q(3)}$ is not coarsening monotonic. For $|\psi\rangle^{AB}|\psi\rangle^{CD}$, $C_{q(2)}(|\psi\rangle^{AB}|\psi\rangle^{CD})=\sqrt{(h_{A}+h_{C})/2}$ may be not larger than $C_{q(2)}(AB|C|D)=\sqrt{2h_{C}/3}$, i.e., it is not tightly coarsening monotonic. Replacing $C_{q(k)}$ with $C_{\alpha(k)}$ in arguments above, we get that $C_{\alpha(k)}$ has the same property as that of $C_{q(k)}$.
Consider $|\psi\rangle^{AB}|\psi\rangle^{CD}$, we have
$C_{G,q(2)}(|\psi\rangle^{AB}|\psi\rangle^{CD})=\sqrt{(h_{A}+h_{C})/2}$
$<$ $C_{G,q(2)}(|\psi\rangle^{AB})+C_{G-q(2)}(|\psi\rangle^{CD})=\sqrt{h_{A}}+\sqrt{h_{C}}$
whenever $h_{A}h_{C}>0$. So it is not additive. Let $|\psi_{1}\rangle=|\psi\rangle^{ABC}|\psi\rangle^{D}$ with $\rho^{A}=\rho^{B}=\rho^{C}$. It turns out that
$C_{G,q(3)}(|\psi_{1}\rangle)=\sqrt[20]{2h_{A}^{10}/9}>C_{G,q(2)}(|\psi_{1}\rangle)=\sqrt{3h_{A}/4}.$
Namely, it is not $k$-monotonic. Let $|\psi_{2}\rangle=|\psi\rangle^{AB}|\psi\rangle^{C}|\psi\rangle^{D}$. Then
$C_{G,q(2)}(|\psi_{2}\rangle)=\sqrt{2h_{A}}/2<C_{G,q(2)}(|\psi\rangle^{AB}|\psi\rangle^{D})$
$=$ $C_{G,q(2)}(A|B|CD)=\sqrt{2h_{A}/3},$
i.e., it is neither coarsening monotonic nor tightly coarsening monotonic. $C_{G,\alpha(k)}$ has the same property as that of $C_{G,q(k)}$.
By definitions, both $E_{(k)}$ and $E^{\prime}_{(k)}$ are not genuine $k$-partite entanglement measures.
5 Examples
[…]